A basketball player is fouled while attempting to make a basket and receives two free throws. The opposing coach believes there is a 34% chance that the player will miss both shots, a 22% chance that he will make one of the shots, and a 44% chance that he will make both shots. a. Construct the appropriate probability distribution. (Round your answers to 2 decimal places.) xPx x) b. What is the probability that he makes no more than one of the shots? (Round your answer to 2 decimal places.) Probability c. What is the probability that he makes at least one of the shots? (Round your answer to 2 decimal places.) Probability
Answer #1

a)

x

P(X=x)

0

0.34

1

0.22

2

0.44

b) Probability that he makes no more than one of the shots=

                  = P(X=0) + P(X=1)

                 = 0.34 + 0.22

               =0.56

c) Probability that he makes at least one of the shots=

                          = P(X=1) + P(X=2)

                         = 0.22 + 0.44

                          =0.66